if the ratio of PVm and RT for a real gas is x/24 at a temp where dP/dVm is 0 find value of 10x

Grade:12th pass

1 Answers

SANJAY SHEKHAWAT
6 years ago
 dp/dvm=0 means critical conditonn in which t=8a/27Rb  v=3b p=a/27b*2  put this value in pv/Rt=x/27  and you will get ans 9  but you asked 10x=90